RFBG Commissioning with Adaptive Secondary - 31 March 2010 UT

Night 12:

Observers: JHeidt, JHill (LBTO)
Telescope Operator: JHill (until 9:30) then WWack
Software: TEdgin
Instrument Support:: JMorris
Telescope Support: JLittle
Adaptive Support: RBriguglio (LBTO)

Summary

Weather was clear through most of the night, strong wind (always > 10 m/s), seeing from 1" (begin) over 0.8" (midnight) to 2" (at the end)

Lost about 2.5 hours due to shell issues.

Tested the latest GCS patch at RFBG. It looks like there could be a reversal of the fixed field aberrations, although convergence is OK. This becomes Issue 2647.

Checked Doug's fixed field aberration corrections, did field aberration data without tip/tilt.

Did some transformation data at bad seing.

Checked x,y-position of probe relative to focus-position.

Preparations

Single frame names refer to: irtc.20100331.xxxxxx.fits

Cube names refer to: irtc.20100331_00xxx_cube.fit

Doug's comments on the problems from last night

Point 3) can be solved with changes in two
parameters (phhh, no new parameters needed!).

The file on the mountain

/lbt/tcs/current/etc/GCS/AIP_IRTC_R.cfg

The parameter that need to be changed are

WFS_rotAngleOffs        0.0             double  # rotator angle offset
between rotator and AGw camera if probe in center

changed from 0.0 to 180.0

The file on the mountain

/lbt/tcs/current/etc/GCS/AIP_IRTC_R_RM_sub11.cfg

The parameters that need to be changed are
and the signs changed on z7, z8, z9, z10

scale_Z7        -1.40   double
scale_Z8        1.40    double
scale_Z9        1.70    double
scale_Z10       -1.70   double

Should become

scale_Z7         1.40   double
scale_Z8        -1.40    double
scale_Z9        -1.70    double
scale_Z10       1.70   double

I think it is best to start the night with the current parameters and
then, once the telescope is collimated, change these parameters
and try to collimate again.  Also, collimate off-axis.

We should see a bit of difference in an IRTC image taken with
a guide start far off-axis.  When we take another set of AGW
Transform data, I will take a look

Andrew will be able to check this new setup next time we
take a set of field aberration data.

**Thus, Torsten, we do not need a new build of GCS
today.

Plan for the night (seeing dependent): Work on collimation issues, collect field aberration measurements and some more transformation data

Details

These details are from Jochen's perspective as the observer. John's notes as the operator are appended below.

Open

01:40 Open, sky mostly clear, few cirrus in northwest, T_out = 5.C, T_Mirror = 3.2C, Hum = 25%, windy

Initial Pointing and Collimation Adjustment

Each night in this RFBG commissioning run, we go to a bright star (Pollux) at Elevation 80 degrees. Then we adjust the vertex tip-tilt (RX/RY) of the secondary to get rid of the coma that we see in the extrafocal pupil on the IRTC.

01:48 pointto Pollux gs=0, TelM = track, PAMode=no, no star, adjust to yesterday values (CA=-77, IE=-48), EL = 78, do ptspiral, OK, star there, adjust CA=-60, IE=-78

Shape of the star looks funny, shell not ready, took a little while. Remember to keep the M1 collimation updated is cases like this where there are delays.

Wind ~ 14 m/s

DX M2 Global X = 2.5, Y = -3.63, Z = 1.77, RX = -360, RY = -120 , coma, adjust to X = 2.5, Y = -3.36, Z = 1.77, RX = -600, RY = 40, CA=-30, IE=-32*

02:47 pointto M5_1210 gs=0, TelM = track , EL = 27, AZ = 170, at very-low elevation to adjust AZ-zeropoint, pointing OK, star found, Adjust IA = -420, keep IE=-32, CA = -30, changed Z = 2.35 to be a bit better in focus

Shape of the star looks again funny, shell need to be reset again, but troubles! No observations with the right hand side of the telescope possible for now.......

Wind ~ 14 m/s

Do some tests with new patches of GCS at left side

Authorize left side for IRTC, register IRTC in IRAF, (in "split" mode)

04:39 pointto Pollux gs=0, TelM = acquire, PAMode=no (preset a couple of times), succeeded in the end

04:50 pointto BS9143 gs=17 TelM=active PA=INDEF=168.97 - collimates! (on the Y-axis) After the fact we see that GCSL sent -860 Z5 in the first two iterations. This was the first left-side collimation, so this initial Z5 may be real.
Mar 31 04:53:47 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
Mar 31 04:54:32 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
Mar 31 04:55:15 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 -661.123 -26.3639 0 0 -14.4446 25.3992 18.838 0 0 0 0 0 0 0 0 0 0 0
Mar 31 04:56:02 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 -201.01 -137.543 0 0 0.342752 -8.58413 -39.3757 0 0 0 0 0 0 0 0 0 0 0
Mar 31 04:56:46 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 10.4665 -28.6949 0 0 1.36672 -21.0556 -4.30327 -0 0 0 -0 -0 0 0 -0 -0 0 -23.5351
Mar 31 04:57:34 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 31.1608 7.88432 0 0 -8.793 4.11296 5.3062 -0 0 0 -0 0 0 0 -0 -0 -0 -20.0981
Mar 31 04:58:20 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 -128.154 -71.3995 0 0 12.1215 -2.49659 -18.6972 -0 0 0 -0 0 0 0 -0 -0 -0 -16.9526
Mar 31 04:59:08 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 33.6407 76.5744 0 0 12.3026 4.56205 3.74787 -0 0 0 -0 0 0 0 -0 -0 -0 -15.6018
Mar 31 04:59:55 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 60.4021 47.3483 0 0 -11.4741 4.26303 10.5529 -0 -0 0 -0 0 0 0 0 -0 -0 -17.6281
Mar 31 05:00:40 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 27.0101 32.867 0 0 -13.8337 -16.4083 -2.50739 -0 0 0 -0 0 0 0 0 -0 -0 -13.0044

05:01 pointto BS9143 gs=17 TelM=active PA=128.0 (to move off the Y-axis) - collimates. NOTE: The off-Y-axis star appears elongated when we first arrive there. GCSL sent -600 Z6 to correct.
Mar 31 05:02:33 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 81.8595 -600.136 0 0 -4.72605 8.77275 37.6049 0 0 0 0 0 0 0 0 0 0 0 Secondary: z(04) = 43.7521 z(07) = 171.5718 z(08) = 131.1049
Mar 31 05:03:18 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 69.5762 -4.72856 0 0 4.08287 -14.0203 1.31671 0 0 -0 -0 0 0 0 -0 -0 0 -5.49523
Mar 31 05:04:04 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 5.75181 -154.076 0 0 33.6704 20.5063 -14.3815 0 0 -0 -0 0 0 0 -0 -0 0 0.330933
Mar 31 05:04:52 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 72.3476 142.025 0 0 -62.2072 25.858 10.8865 0 0 -0 -0 0 0 0 -0 -0 0 -6.44393
Mar 31 05:05:37 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 -64.3001 -41.7626 0 0 11.4168 21.6475 -7.54708 0 0 -0 -0 0 -0 0 -0 -0 0 -7.12027
Mar 31 05:06:22 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 -8.02054 19.8932 0 0 -3.9185 -63.7235 12.7773 0 0 -0 -0 0 0 -0 -0 -0 0 -8.26777
Mar 31 05:07:06 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 22.4249 17.4596 0 0 1.75483 24.9171 0.537334 0 0 -0 -0 0 -0 0 -0 -0 0 -4.79485

05:07 Z9=+300 to SX M1. It converges without Z10 crosstalk, so rotation angles off-axis are good.
Mar 31 05:07:24 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 0 0 0 0 300 0 0 0 0 0 0 0 0 0 0 0 0 0 (manual disturbance to Z9)
Mar 31 05:07:52 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 17.1192 4.60837 0 0 -203.507 -8.66868 -4.36195 0 0 0 0 0 0 0 0 0 0 0
Mar 31 05:08:37 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 15.0721 85.0339 0 0 -76.4608 16.1084 -1.00718 0 0 -0 -0 0 0 0 -0 -0 -0 -5.92257
Mar 31 05:09:26 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 85.3128 -13.4775 0 0 -2.26975 -13.8427 -0.727909 0 0 -0 -0 0 0 -0 -0 -0 0 -2.82519
Mar 31 05:10:12 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 -178.542 -86.4552 0 0 5.51371 23.1294 -17.0522 0 0 0 0 0 0 0 0 0 0 0
Mar 31 05:10:56 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 48.371 -7.07337 0 0 -38.9906 3.95663 6.87231 0 0 -0 -0 0 -0 0 -0 -0 0 -6.36842

05:11 (on-axis) pointto BS9143 gs=0 TelM=active PA=INDEF=0 - Star was distorted in acq image, so we suspect fixed field aberrations may be wrong. GCSL sends negative Z6 to correct.
Mar 31 05:12:37 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0 Secondary: z(04) = 913.8580 z(07) = -175.6008  z(08) = -144.1758
Mar 31 05:13:20 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 -180.531 -250.196 0 0 9.05996 -51.9003 17.3465 0 0 0 0 0 0 0 0 0 0 0
Mar 31 05:14:02 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 26.3148 -152.752 0 0 42.0815 39.8958 -7.60951 -0 -0 -0 0 -0 -0 0 -0 -0 -0 -10.1538
Mar 31 05:14:48 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 -50.1693 -39.3495 0 0 0.0569338 6.15298 5.60756 -0 -0 -0 0 0 -0 0 0 -0 -0 -6.36766
Mar 31 05:15:33 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 -63.8137 92.7809 0 0 30.7282 -40.174 -29.3865 0 0 0 0 0 0 -0 0 -0 -0 -1.17845
Mar 31 05:16:21 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 72.5207 42.5655 0 0 1.52887 9.6366 37.982 0 -0 -0 0 -0 -0 0 0 -0 -0 -9.31275
Mar 31 05:17:08 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 -47.3962 173.29 0 0 -29.7078 -20.7703 -0.706883 0 0 0 0 0 0 0 0 0 0 0
Mar 31 05:17:54 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 -8.34571 -139.334 0 0 -1.69129 -3.73795 -17.3024 -0 0 -0 0 0 -0 0 0 -0 -0 -4.31587
Mar 31 05:18:37 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 28.4943 4.09355 0 0 -0.79832 8.54575 -6.52309 -0 -0 0 0 0 0 0 0 -0 -0 -3.77776
Mar 31 05:19:24 tcs1 LBT_PSFL: ACTOPT: Primary Zernikes 0 0 0 65.0209 75.6006 0 0 21.0088 -24.9471 -0.392846 0 -0 -0 0 -0 0 0 -0 -0 -0 -4.11237

=> Converges on- and off-axis, but apparently has reversed fixed field aberrations as star is elongated when you come back on axis after collimating off-axis. Unfortunately, split mode means that coma terms (the ones we most want to see) are going to the secondary hexapod, and don't get logged in the syslog.

Shell works now, go back at right side

Checked Doug's fixed field aberration corrections

05:25 pointto M5_0347 gs=0, TelM=track, no star, ptspiral, were facing into the wind, go to another M5-star towards east

05:32 pointto M5_0531 gs=0, TelM=track, OK, CA=2, IE=-28, correct for COMA, RX=-460

05:38 pointto M5_0531 gs=0, TelM=active, OK

05:40 pointto BS9138 gs=0, TelM=active, OK, too low elevation (EL=40)

05:40 pointto BS9148 gs=0, TelM=active, EL=76, OK, collimate, OK, guider reports ~ 0.8", its now clear outside

05:48 pointto BS9148 gs=4, TelM=active, star not found

05:55 pointto BS9148 gs=0, TelM=acquire, check pointing, OK, CA=-15, IE=-55

05:58 pointto BS9148 gs=4, TelM=active, OK, converges

06:09 pointto BS9148 gs=4, TelM=active, PA=140,

06:16 => JHill changed WFS_rotAngleOffs from 0 to 180 (based on Dougs recommendation), reversed sign on Z7, Z8, Z9 and Z10, restart GCS

06:21 pointto BS9148 gs=4, TelM=active, PA=140, primary mirror collimation reached limits (no idea where this message comes from, collimation still OK!)

Field aberration data without tip/tilt on BS9152

Data are in /OldRepository/AGW_Data/100331

JHill has modifed his script, starts with 5 images (agw_collimate) on-axis and does always 3 images on-axis afterwards.

06:30 pointo BS9152 gs=0, TelM=active,

06:40 pointto BS9152, GS=0, TelM=acquire, PAM=no, center star, IE=-51, CA=-4

06:45 pointto BS9152 gs=0, TelM=active,

08:40 pointto BS9152 gs=0, TelM=guide, OK centered

M1: Total collimation: *X=-0.828, Y=1.621,Z=0.088, RX=3.663, RY=16.442, RZ=0*
Global offsets: *X-1.331, Y=0.193, Z=1.0, RX=-11.6, RY=2.5, RZ=0*

M2: Total collimation: *X=2.5, Y=-3.63, Z=2.0, RX=-460, RY=-40, RZ=0*
Global offsets: *X=2.5, Y=-3.63, Z=2.0, RX=-460, RY=40, RZ=0*

06:57 cl > newfield_BS9152_right.cl, seeing is variable

GStar wfscimage#
Sequence 1
0 1 - 5
17 6 - 8
0 9 - 11
18 12 - 14
0 15 - 17
7 18 - 20
0 21 - 23
4 24 - 26
0 27 - 29
1 30 - 32
0 33 - 35
9 36 - 38
0 39 - 41

07:45 pointto BS9152, GS=0, TelM=acquire, PAM=no, center star, IE=-50, CA=-15

Shell needs to be reset

07:52 pointto BS9152, GS=0, TelM=active, to collimate in the meantime

07:54 cl > newfield_BS9152_right.cl,

GStar wfscimage#
Sequence 2
0 42 - 46
17 47 - 49
0 50 - 52
18 53 - 55
0 56 - 58

Stopped, enough data to check the fixed aberrations.

At this time we had doubts about whether Doug had updated the coefficients in IDL. He confirmed the following day that he had indeed changed them and we were using here the same updated numbers as we changed for GCS above.

Transformation data

08:14 pointto StoneLagw, GS=0, TelM=active, star at the edge

08:18 pointto StoneLagw, GS=0, TelM=acquire, PAM-, CA=0, IE=-53

08:21 pointto StoneLagw, GS=0, TelM=active, collimate, guider reports ~ 1.2"

08:28 cl < tran_StoneL_mix4.cl,

forgot to turn off movie-mode at IRTC

08:53 pointto StoneLagw, GS=0, TelM=active, collimate, guider reports ~ 1.2"

08:55 cl < tran_StoneL_mix4.cl,

PA = 0

transtar (24) 085659 cube 136
transtar (30) 085937 cube 137
transtar (36) 090211 cube 138
transtar (42) 090451 cube 139
transtar (35) 090732 cube 140

PA = 210

transtar (28) 091035 cube 141
transtar (8) 091241 cube 142, preset failed, probe position was out of range
transtar (7) 091518 cube 143
transtar (3) 091801 cube 144

PA = 315

transtar (6) 092058 cube 145
transtar (5) 092332 cube 146
transtar (24) 092620 cube 147
transtar (30) 092858 cube 148
transtar (36) 093135 cube 149
transtar (12) 093340 cube 150, preset failed, probe position was out of range

PA = 150

transtar (40) 093642 cube 151
transtar (28) 093923 cube 152
transtar (44) 094207 cube 153
transtar (46) 094446 cube 154

PA = 180

transtar (40) 094722 cube 155
transtar (28) 094927 cube 156 preset failed, probe position was out of range
transtar (44) 095127 cube 157, preset failed, probe position was out of range
transtar (7) 095412 cube 158

Seeing gets worse all the time, partly over 2" at the end

09:55 pointto StoneLagw, GS=0, TelM=active, colllimate, stopped (possibly jumps due to the wind)

10:08 pointto StoneMagw, gs=0, TelM=active, collimate,

10:18 cl < tran_StoneM_mix1.cl,

PA = 0

transtar (39) 101658 cube 159
transtar (72) 101946 cube 160
transtar (52) 102226 cube 161
transtar (43) 102503 cube 162
transtar (35) 102738 cube 163
transtar (13) 103024 cube 164

PA = 90

transtar (55) 103304 cube 165
transtar (64) 103542 cube 166
transtar (58) 103820 cube 167
transtar (85) 104106 cube 168
transtar (79) 104350 cube 169
transtar (74) 104626 cube 170
transtar (63) 104909 cube 171
transtar (49) 105151 cube 172
transtar (55) 105432 cube 173

PA = 180

transtar (55) 105728 cube 174
transtar (76) 110011 cube 175
transtar (78) 110251 cube 176
transtar (60) 110535 cube 177
transtar (54) 110815 cube 178

Stopped sufficient data points taken at bad seeing conditions

Photometric ZP's for guider (not useful!)

11:14 pointto M92astrom, GS=0, TelM=active, collimate, 2 stars,

11:25 pointto M92astrom, GS=0, TelM=track PAM-, center on the brighter star ( CA=-13, IE=-55 )

In the meantime cirrus coming closer, data probably not useful

11:31 guider images, 3 sec each, r', 2-4, V, 5-7, B, 8-10

offset: CA=-8, IE=-45, to get the fainter star in the field M92astrom in center of guider

11:40 guider images, 3 sec each, r', 11-13, V, 14-16, B, 17-19

11:43 pointto M92astrom, GS=0, TelM=acquire PAM-, center on the brighter star ( CA=-13, IE=-55 ), OK,

all guider images are rubbish here

Check x,y-position of probe relative to focus-position

11:57 pointto BS9178 gs=0, TelM=active, out of field

12:02 pointto BS9178 gs=0, TelM=acquire, PAM-, CA=7, IE=-40

12:05 pointto BS9178 gs=0, TelM=active, OK, collimate

14:09 turn of guiding, do a set of 100 images with changing focus

focus 21 Guiderimages 21- 35( 21-25 are weird)
focus 31 Guiderimages 36 - 47, yes, out of focus
focus 11 Guiderimages 48 - 59, yes, out of focus
focus 21 Guiderimages 60 - 71
accuracy of the above table +- 1 image

Adjust IAA for IRTC orientation

12:24 pointto BS9178 gs=0, TelM=active, failed, restart PCS

12:28 pointto BS9178 gs=165, TelM=active, OK

doesn't work since Dave's program doesn't account for IAA-values and selecting PA's??? (or something like that)

Close

12:45 Closed, still cirrus, T_out = -0.3., T_Mirror = 0.3, Hum = 30%

John's Night Log as telescope operator for 20100331 UT

JHill acting as operator for first half of night, as Wayne just got back from adventures in Safford, and hasn't had any sleep yet.

BCurtis did walk-thru. JMorris filled LBCs.

Remote regulation was removed from DX primary mirror cell power supplies.

01:08 Stop and restart PCSGUI (was 2880m) and PresetGUI for memory leak control. 01:10 DAshby finishing up Azimuth friction tests after JHoward tuned AZ Pad #4. Azimuth friction is significantly improved.

01:10 Can't talk to agw-control. It is pingable, but ssh command hangs.

01:14 elready, retract stow pins (az and rotator were already running)

Can't talk to lbtmu105 aka agw-control. Called JDavis. Connect with KVM switch. Seems to be out of swap space "Free Swap 0 kB" is repeating along with many other messages. Press power switch to reboot lbtmu105

01:26 azloadfromabs elloadfromabs to reset encoders at 120, 90.

01:28 Az to 90, El to 20

01:30 Open side vent doors.

Succeed in contacting the rebooted agw-control.

01:32 stopAGW -u 2

01:33 stop both GCSL and GCSR

01:34 startAGW -u 2

01:35 home the AGW motors, start GCSR, GCSL

01:39 Clear all active optics on right and initialize lookup tables.

01:42 start opening main shutter doors

Open

Sky is photometric. T=+4.8 degC, D=-13 degC, Wind is moderate (5-10 m/sec).

01:47 Doors are open, start mirror ventilation on both sides.

01:49 Preset to Pollux, spiral search to find star. CA=-60, IE=-78

01:52 Runa is finishing the flattening of the shell. He and Armando had to exclude an actuator from the control loop.

02:09 Start AOSR.

02:16 turn on LFBG rotator (just to verify it is working OK) then turn off.

02:17 startAGW -u 3, but it wouldn't home. So stopAGW -u 3

02:19 startAGW -u 3, then home AGW3

02:22 Stop then Start GCSL.

02:28 Collimate left side of telescope with PSFL

02:31 AdSec is ready after they tracked down a wrong configuration file that was installed today.

02:34 Release brake, Init, and Home Right M2 hexapod.

02:37 Shell RIP'd

02:40 All vent doors to midway.

02:45 Jochen tilts M2 to adjust coma. Need to update M1 collimation before doing this. IE=-32 CA=-30 RX -600 RY +40 Still a bit of Z6.

02:46 Close all vent doors as wind is increasing.

02:49 Preset to low elevation star. Adjust IA=-420 (previously -411).

02:50 Shell has RIP'd again.

02:57 Wind is 8-14 m/sec from the West. Fairly smooth, not especially gusty. We are working cross-wind to the South.

03:00 LWT of winter chiller is -20.3 degC. LWT of Aux Chiller is +7.2 degC.

Instrument cooling is 3.68 degC. Chamber temp is +4.5 degC.

03:05 HBS reservoir is 5.6 degC, and the EL pads are +12 degC (about 6 degC too warm) JLittle goes up to check the circulation pumps for the oil cooling. He found that one circulation pump was off (#2), and that the 10 micron filter on pump #1 was red. So we weren't getting sufficient oil cooling.

03:42 AOSR restarted.

03:47 Home Right M2 Hexapod again after another M2 power cycle.

Left Side Observations

Experiment to be sure that left side Active Optics are still working. (while waiting to get the shell running)

04:03 Stop PCS and change from left to right. Restart PCS Set IA to -411.

04:04 Jochen presets left side to Pollux. First time failed because I forgot to turn on the lfbg rotator.

04:11 Experiment ends for lack of M3 mirror in beam on left side.

04:16 slew to zenith for another try at setting the shell there. Armando and Runa report that they are unable to set the shell because it is shutting down because of some internal diagnostic.

04:20 Oil reservoir temperature is down to -0.2 degC, but winter chiller has been struggling to get it there as LWT is up to -15.5 degC.

04:30 JLittle deploys M3 on left side and rebalances telescope while we are at zenith. Left M3 cables are not connected, so we require simulation mode.

04:37 edit lbt.conf to make M3 in simulation mode

04:39 Left side preset again. failed

04:40 stop and restart OSS (to read the change to lbt.conf)

04:41 Preset succeeds

We find Pollux at these coords on Left. (and we saw it pass by on the IRTC on the right, so we are not so far from co-pointing)

*** Pointing Model Terms *** LEFT
IA      -411.00000
IE      -126.00000
NPAE       6.04950
AW        -3.43410
AN        22.53590
CA        -2.00000
TX        13.15460

04:50 Runa succeeds in setting the shell! So our left side experiments will be short.

04:51 Preset ACTIVE to BS star on left side (star on Y-axis) at PA=168. Collimation succeeds. Seeing is varying 0.7 to 1.1 arcsec.

05:00 PA=128 to move star off the y-axis. NOTE: The off-axis star appears elongated when we first arrive there. It sent -600 Z5.

05:07 Send 300 nm Z9 to SX M1. Rotation angles off-axis are good.

05:11 preset on-axis - star was distorted in acq image, so we suspect fixed field aberrations may be wrong.

05:12 big jump of star image (caused by ??? wind), but guider recovered.

05:14 GCS sent -400 Z5 to fix the image.

Wind gusts on balcony are 12 m/sec max.

05:19 stopGuiding on left

BACK TO THE RIGHT SIDE

05:20 stop PCS, change to right, start PCS

OSS remains with M3 (both of them) in SIMULATION mode

05:22 Stop lfbg rotator

05:24 preset to M5 star on RIGHT side, but shell RIP'd just as we got the image. IE=-32 CA=0

05:30 Imaging is bouncing as we are looking in to the wind (WSW), so we turn to an M5 star out of the wind in the east. Runa recovers the shell quickly. IE=-55 CA=+2 RX=-460 RY+40

05:39 Preset ACTIVE to BS star on-axis, but EL only 41. The active optics did correct focus while we were there.

05:44 Preset ACTIVE to another BS Star - have loads of astigmatism because AOSR still thinks we are at 45 degrees rather than 75 degrees. Active Optics took a few iterations, but did a respectable job of fighting this large

05:48 Restart AOArbitrator and then AOSR. Now the arbitrator knows the correct elevation, but it didn't change the shape of the shell.

05:53 Preset off-axis ACTIVE, missed the star and picked up a ghost which is hard to guide on.

05:56 Preset ACQUIRE on-axis - Adjust CA=-14

05:58 Preset ACTIVE off-axis WFS saturated at 10.8 mag in 30 sec. Seeing varies 0.8 to 1 arcsec.

Collimation seems OK, although we saw up to 200 nm in Z9 bouncing around. This is at X=0 Y=112. These are 8 sec WFS exposures, so their may be atmospheric noise.

06:08 Preset ACTIVE off the Y-axis X=72 Y=86

06:10 Runa restarts the process that sends astigmatism to the shell. But it does not seem to be updating the shape.

06:18 Change Doug's parameters WFS_rotAngleOffs from 0 to 180 deg. And changing the sign of Z7-Z10.

06:20 Restart GCSR to be sure they all took effect. Some had already given us some coma, so perhaps restarting was not needed.

06:21 Preset ACTIVE off the Y-axis X=72 Y=86 EL=81

Primary collimation reached limit. But didn't stay there was this a spurious message.

Move to BS9152. Successfully collimated out a load of astigmatism there (the change from 81 to 60 EL)

06:39 Preset ACQUIRE to tune up pointing IE=-51 CA=-4

06:45 preset ACTIVE on-axis Seeing up to 1.2 arcsec

06:54 play the field aberration script - sort of for practice since seeing isn't so good. Note that the Z6 correction seems not to be running on M2, so we expect a secular drift in Z6 with elevation.

07:10 Midnight walkaround reveals split seam in the DX Mirror Ventilation Duct (Issue 2644), and a stray LBC cover laying in the middle of the floor on level 5. Hopefully the cover was poorly stowed on the floor and didn't fall off of LBC. I stowed the cover on Level 5 against the right wall.

07:16 open rear vent doors midway as wind seems to be slowing down. 07:25 HBS reservoir is down to -1.4 degrees.EL pads are down to 5 degC, although this is still about 1-1.5 degC too warm. LWT of winter chiller is holding steady at -15 degC.

07:38 Primary mirrors are 2.8 degC, while Ref Temp is 2.5, and outside ambient is 2.0. (cooling fairly quickly)

07:43 Preset BS9152 ACQUIRE IE=-50 CA=-15

07:48 Pause observing to restart Shell (various people on Skype to correct previous issues) Image is about 4x8 arcsec when shell is RIP'd.

07:51 preset BS9152 on-axis Seeing improved a bit to 0.8 - 1 arcsec.

07:55 Play newfield_BS9152_right.cl for the second round of data

08:16 Move to StoneL for some transformation data.

Apparently astigmatism correction still isn't working correctly.

08:17 ACQUIRE IE=-53 CA=0 M2 Zglobal 2.2

08:22 collimate on-axis at EL=44.

08:28 start transformation data EL=44

09:12 See nice pupil / lenslet illumination fpr X=-103 Y=164. What makes the outer circular stop in this image? And are we running into it (optically) at large negative X?

09:18 What is vignetting this WFS image? X=130 Y=153 wfscimage000189.fits

09:30 Over to Wayne as operator for the remainder of the night.

-- JochenHeidt - 31 Mar 2010
Topic attachments
I Attachment Action Size Date Who Comment
newfield_BS9152_right.clcl newfield_BS9152_right.cl manage 2 K 01 Apr 2010 - 07:19 JohnHill Script to collect field aberration data
Topic revision: r8 - 05 Apr 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback